Q. If vector A=i+j and vector B=i-k, the angle between vector A and B is


A)zero B)180degree C)60degree D)90degree

Grade:11

2 Answers

Pavan kumar
13 years ago

|A| = |B| = 2^1/2

A.B = 1

Hence angle between A and B is cos-1(A.B/(|A||B|)) = cos-1(1/2) = 60 degrees.

pavan shete
13 years ago
DEAR, A=(1,1,0) MAGNITUDE OF A =[A]=SQRT(2); B=(1,0,-1) ;[B]=SQRT(2); A.B=(1*1)+(1*0)+(0*-1)=1 A.B=[A][B]COS(THETA); 1=2*COS(THETA); COS(THETA)=1/2; HENCE THETA =60 DEGREES; OPTION C IS ANSWER.